OpenStudy (anonymous):

A car is traveling at a constant velocity. After two hours the car has traveled 110 miles. After four hours the car has traveled 220 miles. Assuming a linear equation fits this situation, find the distance that the car will travel after five hours

myininaya (myininaya):

hey my screen is normal now s=d/t right? so we have the speed is constant so we can see that s=110/2=55 and s=220/4=55 so the speed is always 55 so we want to know d when t=5 so 55=d/5 multiply 5 on both sides and you will win! :)